Solve the following DE y (4) + 5y + 4y = sin x + cos2x

Please Solve As soon as
Solve quickly I get you thumbs up directly
Thank's
Abdul-Rahim Taysir
Answer #1

Auxilary + Sy2 + 4y - Sinat what (04 + sp’+4) y = sinn + cos2x equation 04+50²+4=0 aDY + 40² + D²+4=0 B (p² + 4) + 1 (D²+4)=0